# ControlAudio: Tackling Text-Guided, Timing-Indicated and Intelligible Audio Generation via Progressive Diffusion Modeling

### v1

@@ 1 Introduction
Text-to-audio (TTA) generation systems aim at synthesizing high-fidelity audio samples that are consistent with the given natural language description, e.g., "A bird is chirping" Liu et al. (2023); Ghosal et al. (2023); Huang et al. (2023); Evans et al. (2024a). Recent efforts are exploring more fine-grained control for TTA systems, which can be categorized into two main classifications. The first group adds precise timing control, e.g., "A bird is chirping, at 2-5 seconds", with innovations spanning conditioning techniques Wang et al. (2025c); Xie et al. (2024) and training-free latent manipulation Jiang et al. (2025). The second group works on intelligible audio generation, e.g., "A bird is chirping, and a man is saying: ’it’s a very sunny day’", by introducing additional modules to encode both audio and speech semantic information Lee et al. (2024b); Jung et al. (2025). However, as expensive to collect large-scale text-audio datasets with precise timing and speech information, their controllable generation performance at scale remains limited, and none of the prior work explores timing-controlled and intelligible TTA generatione.g., "A bird is chirping, at 0-5 seconds, and then a man is saying: ’it’s a very sunny day’, at 7-10 seconds", within a unified framework.
@@
With target distributions predefined by datasets constructed above, we introduce progressive diffusion training, fulfilling high-quality TTA synthesis at pre-training and gradually integrating fine-grained control signals at continual learning stages. At the first stage, we pre-train a diffusion transformer (DiT) in the latent space directly compressed from the waveform space, solely conditioned on text indication, achieving high-fidelity TTA at scale. At the second stage, we fine-tune the latent DiT on both text and timing conditions, enabling the model to precisely control the timing windows of each sound event. In controllable TTA generation, a common issue is the sacrifice of text-conditioned synthesis quality without fine-grained conditions Wang et al. (2025c). Hence, in ControlAudio, we switch the condition between the text condition and the $\langle\text{text, timing}\rangle$ condition at the second stage, avoiding catastrophic forgetting in progressive training. At the final stage, given the audio generation prior learned in prior stages, we continually train the diffusion model by switching the condition among text, $\langle\text{text, timing}\rangle$, and $\langle\text{text, timing, phoneme}\rangle$, achieving high-fidelity audio synthesis conditioned on flexible indication.
@@ 2.1 Controllable TTA Generation
Recent works have aimed to add temporal control to TTA models, primarily through two strategies. Training-based methods, such as MC-Diffusion Guo et al. (2024) and PicoAudio Xie et al. (2024), condition on predefined event classes, limiting their expressiveness for open-domain prompts. While AudioComposer Wang et al. (2025c) uses natural language, it struggles with ambiguity in complex event descriptions. Conversely, training-free approaches like TG-Diff Du et al. (2024) and FreeAudio Jiang et al. (2025) enforce alignment during inference, but often incur high computational costs and fail in dense scenarios. A parallel challenge is the generation of intelligible speech. Most existing TTA models render speech as vague vocalizations. While models like VoiceLDM Lee et al. (2024b) and VoiceDiT Jung et al. (2025) can synthesize high-quality speech in context, they operate as specialized TTS systems and lack control over general audio events. Furthermore, prior work, including specialized models for controllable dialogue like CoVoMix2 Zhang et al. (2025), has largely focused on single-speaker or speech-only scenarios. This work is the first to address these dual challenges, proposing a unified framework for the timing-controlled, joint generation of both general audio events and intelligible multi-speaker dialogue.
@@ 2.2 Progressive Modeling
In recent cross-modal generation tasks, such as video or avatar generation conditioned on diverse control signals Lin et al. (2025); Hu et al. (2025), progressive modeling has proven effective in handling multi-condition video generation. However, its advantages have not been extended to controllable TTA generation, where precise timing control and intelligible speech represent critical requirements but remain unresolved.
@@ 3.1 Diffusion-based TTA Generation
Diffusion-based Peebles and Xie (2023); Li et al. (2024) TTA models are typically trained to learn a conditional reverse of a data-to-noise forward process Ho et al. (2020), progressively removing noise from an initial random state conditioned on a text prompt over multiple diffusion steps. This framework consists of three main modules: 1) an audio varational autoencoder (VAE), responsible for transforming the audio sample into a compressed latent representation while ensuring the reconstruction quality; 2) a pretrained text encoder, which encodes a text prompt into conditioning embeddings; and 3) a latent diffusion model, which predicts the denoised audio latents conditioned on the text embeddings. In ControlAudio, we employ a DiT-based architecture to ensure scalability Evans et al. (2025), conditioned on the text, timing, and phoneme embeddings to generate the latent audio representation directly compressed from the waveform, without cascaded decoding Liu et al. (2023); Xie et al. (2024); Guo et al. (2024).
@@ 3.2 Classifier-Free Guidance
Classifier-Free Guidance (CFG) Ho and Salimans (2022); Wang et al. (2025a) emphasizes the guidance of a conditioning signal $c$ during sampling. At each sampling step, CFG-guided diffusion models produce two predictions: a conditional estimation $\epsilon_{\theta}(x_{t},c)$ and an unconditional estimation $\epsilon_{\theta}(x_{t},\emptyset)$. Then the final prediction is obtained by extrapolating these two terms with a guidance scale $w>1$:
@@ 4.2 Dataset Construction
Annotated Data. Our data annotation pipeline begins with the AudioSet-SL Hershey et al. (2021) dataset, chosen for its reliable temporal annotations while lacking corresponding speech transcripts. To create the ControlAudio dataset, we first select all clips containing "human speech" and then extract a clean speech track from each using a dual-demixing strategy inspired by MTV Weng et al. (2025) that leverages both MVSEP Solovyev and Spleeter Hennequin et al. (2020). The clean track is subsequently segmented into individual events using the original timestamps. Finally, each segmented event is transcribed using Gemini 2.5 Pro [^1] https://deepmind.google/models/gemini/pro/. Further details of this entire pipeline are provided in Appendix A.3. This transcription process enriches the dataset i.e., expanding condition to $\langle\text{text, timing, phoneme}\rangle$ for fine-grained control. For example, a generic annotation like (man speaking, <3.00,5.00>) is transformed into a specific, content-rich event (man speaking: "It’s been raining all day.", <3.00,5.00>).
Simulated Data. To further expand our dataset, we construct a large-scale simulated dataset guided by real-world data distribution. Specifically, we first analyze the AudioSet-SL dataset to derive statistical priors on speech activity patterns, with further details provided in Appendix A.4. These distributions guide our synthesis process, which proportionally simulates two main scenarios: single-speaker scenarios (monologue) are created by combining multiple utterances from the same speaker in LibriTTS-R, while multi-speaker scenarios (dialogue) are formed by sampling from different speakers. After composing the speech samples, we simulate a plausible temporal arrangement for the utterances. Finally, the composed speech is mixed with non-speech backgrounds from WavCaps Mei et al. (2024) and VGG-Sound Chen et al. (2020) at a signal-to-noise ratio sampled from a uniform 2 to 10 dB range Jung et al. (2025). Through this simulation pipeline, we generate an additional 171,246 complex audio scenes, significantly expanding the scale and diversity of our training data.
@@ 4.3 Unified Semantic Modeling
To address the challenge of encoding diverse condition information, including text, timing, and phoneme features, we propose a unified semantic modeling approach that handles them with a single text encoder in a progressive and coarse-to-fine manner. This approach avoids the complexity of multiple specialized modules Lee et al. (2024b) by first establishing a robust structural representation, providing a simple yet effective solution for rendering fine-grained content in audio generation.
@@
Structured Prompt for Phoneme Representation. Our approach to synthesizing intelligible speech is built directly upon the temporal foundation provided by the structured prompt. A key insight of our work is that the explicit timing windows (<start,end>) assigned to each speech event inherently define the utterance’s total duration. This is a significant advantage over standard TTS systems Anastassiou et al. (2024); Lee et al. (2024a), which must employ complex, often error-prone models just to predict the duration of each phoneme or word. By having the duration as a given constraint, our framework can bypass this challenging duration modeling task entirely.
@@ 4.4 Progressive Model Training
To train our model for multi-condition audio generation, we adopt a progressive three-stage training strategy. This approach allows the model to acquire fine-grained control capability incrementally, where each new stage builds upon and refines the skills learned previously, ensuring a stable and efficient learning process. At each stage, the model is optimized using the conditional diffusion objective Ho et al. (2020), where a network is trained to predict the noise $\epsilon$ added on the clean audio latents:
@@
Stage 1: TTA Pre-training. We first pre-train a DiT Evans et al. (2025) on large-scale text-audio datasets to learn a robust, general mapping from textual descriptions to audio latent representation, ensuring high-fidelity text-guided audio generation.
@@ 5.1 Experiment Setting
Evaluation Datasets. To objectively evaluate our method, we utilize several established datasets, each targeting a specific capability. For timing-controllable generation, we use the publicly available test split from AudioCondition Guo et al. (2024), whose fine-grained temporal annotations are ideal for this task. For intelligible speech generation, we use the AC-Filtered Lee et al. (2024b) dataset. For evaluating general TTA performance, we report results on the AudioCaps test set Kim et al. (2019). In addition, we include the LibriTTS-R and LibriSpeech Panayotov et al. (2015) test-clean splits for specific ablation studies.
Evaluation Metrics. We conduct a comprehensive evaluation covering three key aspects: temporal control, audio quality, and speech intelligibility. For temporal control, we follow prior work Guo et al. (2024); Wang et al. (2025c) and report two metrics computed by a sound event detection (SED) system Mesaros et al. (2016): the event-based measures (Eb) and the clip-level macro F1 score (At). For audio quality, we employ a suite of standard metrics, including Fréchet Audio Distance (FAD), Kullback–Leibler (KL) divergence, Fréchet Distance (FD), Inception Score (IS) Liu et al. (2023) and CLAP Wu et al. (2023). For speech intelligibility, we conduct both objective and subjective tests. Objectively, we measure the Word Error Rate (WER) by transcribing generated speech with the Whisper Large-v3 model Radford et al. (2023). Subjectively, we conduct Mean Opinion Score (MOS) tests where 20 participants rate three aspects on a five-point scale: Speech Intelligibility, Overall Quality (OVL), and Relevance to the prompt (REL). Further details are provided in the Appendix C.
@@ 5.2 Main Results
Timing-Controlled Audio Generation. We compare ControlAudio with several state-of-the-art TTA models, including AudioLDM Liu et al. (2023), AudioLDM 2 Liu et al. (2024a), Tango Ghosal et al. (2023), and our in-house implementation of Stable Audio Evans et al. (2024a, b, 2025). We also include models that incorporate explicit temporal conditioning signals, such as MC-Diffusion Guo et al. (2024) and AudioComposer Wang et al. (2025c), as well as the training-free baselines TG-Diff Du et al. (2024) and FreeAudio Jiang et al. (2025). TG-Diff reports both timing and audio quality metrics under a training-free framework but relies on a different sound event detection model Turpault et al. (2019) compared to other baselines. Control-Condition-to-Audio (CCTA) is a baseline variant of MC-Diffusion that uses only control conditions without textual input, while Tango + LControl is an AudioComposer variant built on Tango with language-based temporal control. In terms of efficiency, we measure the real-time factor (RTF) Liu et al. (2024b, c) for all models on a single NVIDIA A800 GPU. As shown in Table 1, ControlAudio achieves competitive or superior temporal alignment compared to existing methods, while significantly improving audio generation quality in both objective and subjective metrics, and does so without introducing additional inference overhead compared to baseline models.
@@ 5.3 Ablation Study
Ablation of Vocabulary Granularity. To determine the optimal vocabulary granularity for intelligible speech, we conduct an ablation study on the LibriTTS-R and LibriSpeech test-clean datasets. We first compare three variants of our model, differentiated by their vocabulary: word-level, sub-word (BPE), and phoneme-level. For broader context, we also report results from the strong VoiceLDM baselines. Evaluation metrics include WER for intelligibility and UTMOS Saeki et al. (2022)(UT-M) for speech naturalness. For WER calculation on LibriSpeech, we adopt a HuBERT-based ASR model Hsu et al. (2021), following prior works Shen et al. (2023). As shown in Table 5, the phoneme-level model consistently and significantly outperforms the other granularities, achieving the lowest WER and highest UTMOS scores. These findings confirm that phonemes provide a more direct representation of spoken content, facilitating a tighter alignment between the prompt and the acoustic output, which ultimately translates into substantially clearer and more intelligible speech.
@@ A.1 Pretraining Datasets and Preprocessing
Table 6 provides a comprehensive summary of all corpora used for pretraining our TTA backbone. To learn a robust mapping between text and audio, we aggregate a diverse mixture of large-scale, publicly available datasets. This includes datasets with descriptive captions, such as the large-scale WavCaps Mei et al. (2024) and the widely-used AudioCaps Kim et al. (2019), as well as corpora with high-level event labels, like the massive AudioSet Gemmeke et al. (2017). This rich combination of data sources, spanning both detailed descriptions and a wide vocabulary of sound classes, allows the model to learn robust and versatile semantic representations.

### v2

@@ 1 Introduction
Text-to-audio (TTA) generation systems aim at synthesizing high-fidelity audio samples that are consistent with the given natural language description, e.g., "A bird is chirping" liu2023audioldm; ghosal2023text; huang2023make; evans2024fast. Recent efforts are exploring more fine-grained control for TTA systems, which can be categorized into two main classifications. The first group adds precise timing control, e.g., "A bird is chirping, at 2-5 seconds", with innovations spanning conditioning techniques wang2025audiocomposer; xie2024picoaudio and training-free latent manipulation jiang2025freeaudio. The second group works on intelligible audio generation, e.g., "A bird is chirping, and a man is saying: ’it’s a very sunny day’", by introducing additional modules to encode both audio and speech semantic information lee2024voiceldm; jung2025voicedit. However, as expensive to collect large-scale text-audio datasets with precise timing and speech information, their controllable generation performance at scale remains limited, and none of the prior work explores timing-controlled and intelligible TTA generatione.g., "A bird is chirping, at 0-5 seconds, and then a man is saying: ’it’s a very sunny day’, at 7-10 seconds", within a unified framework.
@@
With target distributions predefined by datasets constructed above, we introduce progressive diffusion training, fulfilling high-quality TTA synthesis at pre-training and gradually integrating fine-grained control signals at continual learning stages. At the first stage, we pre-train a diffusion transformer (DiT) in the latent space directly compressed from the waveform space, solely conditioned on text indication, achieving high-fidelity TTA at scale. At the second stage, we fine-tune the latent DiT on both text and timing conditions, enabling the model to precisely control the timing windows of each sound event. In controllable TTA generation, a common issue is the sacrifice of text-conditioned synthesis quality without fine-grained conditions wang2025audiocomposer. Hence, in ControlAudio, we switch the condition between the text condition and the $\langle\text{text, timing}\rangle$ condition at the second stage, avoiding catastrophic forgetting in progressive training. At the final stage, given the audio generation prior learned in prior stages, we continually train the diffusion model by switching the condition among text, $\langle\text{text, timing}\rangle$, and $\langle\text{text, timing, phoneme}\rangle$, achieving high-fidelity audio synthesis conditioned on flexible indication.
@@ 2.1 Controllable TTA Generation
Recent works have aimed to add temporal control to TTA models, primarily through two strategies. Training-based methods, such as MC-Diffusion guo2024audio and PicoAudio xie2024picoaudio, condition on predefined event classes, limiting their expressiveness for open-domain prompts. While AudioComposer wang2025audiocomposer uses natural language, it struggles with ambiguity in complex event descriptions. Conversely, training-free approaches like TG-Diff du2024controllable and FreeAudio jiang2025freeaudio enforce alignment during inference, but often incur high computational costs and fail in dense scenarios. A parallel challenge is the generation of intelligible speech. Most existing TTA models render speech as vague vocalizations. While models like VoiceLDM lee2024voiceldm and VoiceDiT jung2025voicedit can synthesize high-quality speech in context, they operate as specialized TTS systems and lack control over general audio events. Furthermore, prior work, including specialized models for controllable dialogue like CoVoMix2 zhang2025covomix2, has largely focused on single-speaker or speech-only scenarios. This work is the first to address these dual challenges, proposing a unified framework for the timing-controlled, joint generation of both general audio events and intelligible multi-speaker dialogue.
@@ 2.2 Progressive Modeling
In recent cross-modal generation tasks, such as video or avatar generation conditioned on diverse control signals OmniHuman; hunyuancustom, progressive modeling has proven effective in handling multi-condition video generation. However, its advantages have not been extended to controllable TTA generation, where precise timing control and intelligible speech represent critical requirements but remain unresolved.
@@ 3.1 Diffusion-based TTA Generation
Diffusion-based peebles2023scalable; li2024qa TTA models are typically trained to learn a conditional reverse of a data-to-noise forward process ho2020denoising, progressively removing noise from an initial random state conditioned on a text prompt over multiple diffusion steps. This framework consists of three main modules: 1) an audio varational autoencoder (VAE), responsible for transforming the audio sample into a compressed latent representation while ensuring the reconstruction quality; 2) a pretrained text encoder, which encodes a text prompt into conditioning embeddings; and 3) a latent diffusion model, which predicts the denoised audio latents conditioned on the text embeddings. In ControlAudio, we employ a DiT-based architecture to ensure scalability evans2025stable, conditioned on the text, timing, and phoneme embeddings to generate the latent audio representation directly compressed from the waveform, without cascaded decoding liu2023audioldm; xie2024picoaudio; guo2024audio.
@@ 3.2 Classifier-Free Guidance
Classifier-Free Guidance (CFG) ho2022classifier; wang2025audiomog emphasizes the guidance of a conditioning signal $c$ during sampling. At each sampling step, CFG-guided diffusion models produce two predictions: a conditional estimation $\epsilon_{\theta}(x_{t},c)$ and an unconditional estimation $\epsilon_{\theta}(x_{t},\emptyset)$. Then the final prediction is obtained by extrapolating these two terms with a guidance scale $w>1$:
@@ 4.2 Dataset Construction
Annotated Data. Our data annotation pipeline begins with the AudioSet-SL hershey2021benefit dataset, chosen for its reliable temporal annotations while lacking corresponding speech transcripts. To create the ControlAudio dataset, we first select all clips containing "human speech" and then extract a clean speech track from each using a dual-demixing strategy inspired by MTV weng2025audio that leverages both MVSEP ZFTurbo23 and Spleeter hennequin2020spleeter. The clean track is subsequently segmented into individual events using the original timestamps. Finally, each segmented event is transcribed using Gemini 2.5 Pro [^1] https://deepmind.google/models/gemini/pro/. Further details of this entire pipeline are provided in Appendix A.3. This transcription process enriches the dataset i.e., expanding condition to $\langle\text{text, timing, phoneme}\rangle$ for fine-grained control. For example, a generic annotation like (man speaking, <3.00,5.00>) is transformed into a specific, content-rich event (man speaking: "It’s been raining all day.", <3.00,5.00>).
Simulated Data. To further expand our dataset, we construct a large-scale simulated dataset guided by real-world data distribution. Specifically, we first analyze the AudioSet-SL dataset to derive statistical priors on speech activity patterns, with further details provided in Appendix A.4. These distributions guide our synthesis process, which proportionally simulates two main scenarios: single-speaker scenarios (monologue) are created by combining multiple utterances from the same speaker in LibriTTS-R, while multi-speaker scenarios (dialogue) are formed by sampling from different speakers. After composing the speech samples, we simulate a plausible temporal arrangement for the utterances. Finally, the composed speech is mixed with non-speech backgrounds from WavCaps mei2024wavcaps and VGG-Sound chen2020vggsound at a signal-to-noise ratio sampled from a uniform 2 to 10 dB range jung2025voicedit. Through this simulation pipeline, we generate an additional 171,246 complex audio scenes, significantly expanding the scale and diversity of our training data.
@@ 4.3 Unified Semantic Modeling
To address the challenge of encoding diverse condition information, including text, timing, and phoneme features, we propose a unified semantic modeling approach that handles them with a single text encoder in a progressive and coarse-to-fine manner. This approach avoids the complexity of multiple specialized modules lee2024voiceldm by first establishing a robust structural representation, providing a simple yet effective solution for rendering fine-grained content in audio generation.
@@
Structured Prompt for Phoneme Representation. Our approach to synthesizing intelligible speech is built directly upon the temporal foundation provided by the structured prompt. A key insight of our work is that the explicit timing windows (<start,end>) assigned to each speech event inherently define the utterance’s total duration. This is a significant advantage over standard TTS systems anastassiou2024seed; lee2024ditto, which must employ complex, often error-prone models just to predict the duration of each phoneme or word. By having the duration as a given constraint, our framework can bypass this challenging duration modeling task entirely.
@@ 4.4 Progressive Model Training
To train our model for multi-condition audio generation, we adopt a progressive three-stage training strategy. This approach allows the model to acquire fine-grained control capability incrementally, where each new stage builds upon and refines the skills learned previously, ensuring a stable and efficient learning process. At each stage, the model is optimized using the conditional diffusion objective ho2020denoising, where a network is trained to predict the noise $\epsilon$ added on the clean audio latents:
@@
Stage 1: TTA Pre-training. We first pre-train a DiT evans2025stable on large-scale text-audio datasets to learn a robust, general mapping from textual descriptions to audio latent representation, ensuring high-fidelity text-guided audio generation.
@@ 5.1 Experiment Setting
Evaluation Datasets. To objectively evaluate our method, we utilize several established datasets, each targeting a specific capability. For timing-controllable generation, we use the publicly available test split from AudioCondition guo2024audio, whose fine-grained temporal annotations are ideal for this task. For intelligible speech generation, we use the AC-Filtered lee2024voiceldm dataset. For evaluating general TTA performance, we report results on the AudioCaps test set kim2019audiocaps. In addition, we include the LibriTTS-R and LibriSpeech panayotov2015librispeech test-clean splits for specific ablation studies.
Evaluation Metrics. We conduct a comprehensive evaluation covering three key aspects: temporal control, audio quality, and speech intelligibility. For temporal control, we follow prior work guo2024audio; wang2025audiocomposer and report two metrics computed by a sound event detection (SED) system mesaros2016metrics: the event-based measures (Eb) and the clip-level macro F1 score (At). For audio quality, we employ a suite of standard metrics, including Fréchet Audio Distance (FAD), Kullback–Leibler (KL) divergence, Fréchet Distance (FD), Inception Score (IS) liu2023audioldm and CLAP wu2023large. For speech intelligibility, we conduct both objective and subjective tests. Objectively, we measure the Word Error Rate (WER) by transcribing generated speech with the Whisper Large-v3 model radford2023robust. Subjectively, we conduct Mean Opinion Score (MOS) tests where 20 participants rate three aspects on a five-point scale: Speech Intelligibility, Overall Quality (OVL), and Relevance to the prompt (REL). Further details are provided in the Appendix C.
@@ 5.2 Main Results
Timing-Controlled Audio Generation. We compare ControlAudio with several state-of-the-art TTA models, including AudioLDM liu2023audioldm, AudioLDM 2 liu2024audioldm, Tango ghosal2023text, and our in-house implementation of Stable Audio evans2024fast; evans2024long; evans2025stable. We also include models that incorporate explicit temporal conditioning signals, such as MC-Diffusion guo2024audio and AudioComposer wang2025audiocomposer, as well as the training-free baselines TG-Diff du2024controllable and FreeAudio jiang2025freeaudio. TG-Diff reports both timing and audio quality metrics under a training-free framework but relies on a different sound event detection model turpault2019sound compared to other baselines. Control-Condition-to-Audio (CCTA) is a baseline variant of MC-Diffusion that uses only control conditions without textual input, while Tango + LControl is an AudioComposer variant built on Tango with language-based temporal control. In terms of efficiency, we measure the real-time factor (RTF) liu2024audiolcm; liu2024flashaudio for all models on a single NVIDIA A800 GPU. As shown in Table 1, ControlAudio achieves competitive or superior temporal alignment compared to existing methods, while significantly improving audio generation quality in both objective and subjective metrics, and does so without introducing additional inference overhead compared to baseline models.
@@ 5.3 Ablation Study
Ablation of Vocabulary Granularity. To determine the optimal vocabulary granularity for intelligible speech, we conduct an ablation study on the LibriTTS-R and LibriSpeech test-clean datasets. We first compare three variants of our model, differentiated by their vocabulary: word-level, sub-word (BPE), and phoneme-level. For broader context, we also report results from the strong VoiceLDM baselines. Evaluation metrics include WER for intelligibility and UTMOS saeki2022utmos(UT-M) for speech naturalness. For WER calculation on LibriSpeech, we adopt a HuBERT-based ASR model hsu2021hubert, following prior works shen2023naturalspeech. As shown in Table 5, the phoneme-level model consistently and significantly outperforms the other granularities, achieving the lowest WER and highest UTMOS scores. These findings confirm that phonemes provide a more direct representation of spoken content, facilitating a tighter alignment between the prompt and the acoustic output, which ultimately translates into substantially clearer and more intelligible speech.
@@ A.1 Pretraining Datasets and Preprocessing
Table 6 provides a comprehensive summary of all corpora used for pretraining our TTA backbone. To learn a robust mapping between text and audio, we aggregate a diverse mixture of large-scale, publicly available datasets. This includes datasets with descriptive captions, such as the large-scale WavCaps mei2024wavcaps and the widely-used AudioCaps kim2019audiocaps, as well as corpora with high-level event labels, like the massive AudioSet gemmeke2017audio. This rich combination of data sources, spanning both detailed descriptions and a wide vocabulary of sound classes, allows the model to learn robust and versatile semantic representations.